# Learning Personalized Decision Support Policies

### v2

@@ Introduction
To construct expertise profiles with realistic values for each $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$, we collect data on user decisions across different users and then calculate individual $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$. The set of participant expertise profiles form a population of decision-makers that we refer to as human-informed synthetic decision-makers. From the estimated $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$ of each human-informed synthetic decision-maker, we can simulate decision-maker behavior. We leverage these decision-makers to evaluate multiple values of $\lambda$ in our computational experiments and to select $\lambda$ values to deploy on unseen decision-makers in our human subject experiments.
To construct human-informed synthetic decision-makers, we recruited 20 participants from Prolific (10 for CIFAR-$3A$ and 10 for MMLU-$2A$). We use the same recruitment scheme as the larger human subject experiment described in the Appendix. We define regions of expertise over class labels for CIFAR-$3A$ and over question topics for MMLU-$2A$, as we expect $r_{\textsc{A}_{i}}(x;h)$ to be roughly constant for $x\in{\mathcal{X}}_{j}$ where ${\mathcal{X}}_{j}$ is defined by a class label or question topic. We showed each participant similar inputs with different forms of support to estimate $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$ for each support $\textsc{A}_{i}$ in each region ${\mathcal{X}}_{j}$. On each trial, each participant is randomly assigned a form of support; trials are approximately balanced by the type of support and grouping (i.e., topic or class). We compute participant accuracy averaged over all trials: 100 for CIFAR-$3A$, 60 for MMLU-$2A$. We denote expertise profiles in subsequent experiments as follows: if there were three regions in the input space, an individual’s expertise profile under support $\textsc{A}_{i}$ would be written as $r_{\textsc{A}_{i}}=[0.7,0.1,0.7]$, meaning the individual incurs a loss of $0.7$ on ${\mathcal{X}}_{1}$, $0.1$ on ${\mathcal{X}}_{2}$, and $0.7$ on ${\mathcal{X}}_{3}$.
@@
  • Approximately Invariant expertise across all the regions under different forms of support, i.e., $r_{\textsc{A}_{1}}({\mathcal{X}}_{j};h)\approx r_{\textsc{A}_{2}}({\mathcal{X}}_{j};h)\approx\cdots\approx r_{\textsc{A}_{k}}({\mathcal{X}}_{j};h),\forall j\in[N]$. In the cost-aware setting, a decision support policy should learn the fixed form of support that corresponds to picking the cheapest form of support.
@@
We use human-informed synthetic decision-makers to evaluate Modiste in both cost-agnostic and cost-aware settings.
Cost-agnostic results.
@@
We investigate how personalized policies compare against offline baselines under each expertise profile (Table 1). We verify that learning decision support policies are not helpful for decision-makers with “invariant” expertise profiles. However, for individuals who fall under the “strictly better” and “varying” profiles, we find at least one personalized policy outperforms offline policies and learns a policy that is closer to the decision-maker’s optimal performance. This is because a personalized policy identifies which form of support is better in each context, compared to fixed offline policies which always show one form of support or to the population variant, which may not provide the correct form of support to each individual. We note that KNN generally outperforms LinUCB, the latter of which can be saddled by its implicit linearity assumption.
Cost-aware results.
When accounting for both performance and cost of support, we aim to learn low-cost policies for each decision-maker that are within $\epsilon=0.05$ of their optimal loss $L_{h}^{\text{opt}}$. In these experiments, we focus on MMLU-$2A$, where individuals naturally have “varying” and “strictly better” expertise profiles, two settings it makes sense to consider both loss and cost objectives (Table 2). We omit the “invariant” profile as the best policy is to simply select the cheapest form of support. For individuals who are “strictly better,” it may not be possible to further reduce the expected loss compared to one of the offline policies; however, we find that personalization can reduce cost while still achieving an expected loss that is within $\epsilon$ of $L_{h}^{\text{opt}}$. For the individuals with “varying” profiles, we find that personalization can improve both performance and cost over many of the offline baselines. We find the population baseline is inconsistent in its ability to identify high-performing policies at low cost, in contrast to the personalized policies learned. We report similar results for multiple datasets in the Appendix.
Selecting hyper-parameters.
In the cost-aware setting, we used the human-informed synthetic decision-makers to learn policies for multiple values of $\lambda$ simultaneously. However, we observe that often only a subset of the learned policies fall within $\epsilon$ of $L_{h}^{\text{opt}}$ (e.g., in Figure 8), underscoring that it may be challenging to consistently select a suitable $\lambda$ in practice. We further study the effect of various parameters, e.g., exploration parameters, KNN parameters, embedding size, and the number of interactions in the Appendix.
Table 3: We report expected loss $L_{h}(\pi)$ and expected cost $c(\pi)$ incurred (lower is better) in the last 10 trials by Prolific participants for each Algorithm and bold the variant with the lowest $L_{h}(\pi)$. We also consider different choices of $\lambda$, where $\lambda=1.0$ corresponds to the cost-agnostic setting and $\lambda\neq 1.0$ corresponds to a cost-aware setting, where the choice of $\lambda$ was selected according to our cost-aware workflow.
Algorithm $L_{h}(\pi)$ $c(\pi)$
H-Only $0.68\pm 0.13$ 00
H-Model $0.50\pm 0.11$ $0.5$
H-Consensus $0.32\pm 0.07$ $0.5$
Population $0.66\pm 0.08$ 00
Modiste-LinUCB ($\lambda=1.0$) $0.22\pm 0.15$ $0.38\pm 0.06$
Modiste-KNN ($\lambda=1.0$) $\mathbf{0.1\pm 0.06}$ $\mathbf{0.44\pm 0.07}$
Modiste-LinUCB ($\lambda=0.85$) $0.5\pm 0.15$ $0.14\pm 0.07$
Modiste-KNN ($\lambda=0.75$) $0.14\pm 0.23$ $0.39\pm 0.06$
Algorithm $L_{h}(\pi)$ $c(\pi)$
H-Only $0.51\pm 0.14$ 00
H-LLM $\mathbf{0.22\pm 0.14}$ $\mathbf{0.1}$
Population $0.33\pm 0.19$ $0.06\pm 0.01$
Modiste-LinUCB ($\lambda=1.0$) $0.25\pm 0.11$ $0.05\pm 0.02$
Modiste-KNN ($\lambda=1.0$) $0.3\pm 0.05$ $0.07\pm 0.02$
Modiste-LinUCB ($\lambda=0.95$) $0.35\pm 0.14$ $0.04\pm 0.01$
Modiste-KNN ($\lambda=0.75$) $0.3\pm 0.11$ $0.07\pm 0.02$
Modiste with real users
@@
To validate whether Modiste can improve decision-maker performance and cost in practice, we run a series of human subject experiments (i.e., ethics-reviewed studies with real human participants). We first introduce the set-up of our user study; additional information can be found in the Appendix. To our knowledge, we are one of the first works that run actual user studies to demonstrate the benefits of learning personalized policies online: see an extensive comparison to prior work in the Appendix.
Figure 4: Snapshots of the learned decision support policies computed at the end of $100$ interactions with randomly sampled users for CIFAR-$3A$. The forms of support are colored in t-SNE embedding space. In the cost-agnostic setting, both LinUCB and KNN generally identify the optimal form of support for a given input. Since the choice of support for images of Birds does not matter given the task set-up, we nicely find that when we introduce cost (i.e., when $\lambda\neq 1$), the learned policies favor the cheaper H-Only arm for images of Birds. All participant plots are included in Figure 12.
@@
We recruit a total of $125$ crowdsourced participants from Prolific to interact with Modiste ($N=45$ and $N=80$ for CIFAR-$3A$ and MMLU-$2A$, respectively). We recruit more participants for MMLU-$2A$, as we expect greater individual differences in regions where support is needed, e.g., some participants may be good at mathematics and struggle in biology, whereas others may excel in biology questions, whereas in CIFAR-$3A$, there is an “optimal” form of support for each stimulus.
@@
Each participant is assigned to only one task. Within a task, participants are randomly assigned to one algorithm variant; an equal number of participants are included per variant (i.e., 10 for MMLU and 5 for CIFAR). Participants are required to reside in the United States and speak English as a first language. Participants are paid at a base rate of \$9/hr, and are told they may be paid an optional bonus up to \$10/hr based on the number of correct responses. We allot 25-30 minutes for the CIFAR task and 30-40 for MMLU, as each MMLU question takes more effort. We applied the bonus to all participants in all studies.
Trade-off parameter selection.
For our cost-agnostic experiments, we let $\lambda=1.0$ for each variant across all datasets. For our cost-aware experiments, we follow the strategies specified in the previous section to identify $\lambda$ for each variant. We visualize this process in detail in Figure 9. For CIFAR-$3A$, the selected parameter values were $\lambda=0.85$ for LinUCB and $\lambda=0.75$ for KNN. For MMLU-$2A$, the selected parameter values were $\lambda=0.95$ for LinUCB and $\lambda=0.75$ for KNN.
Assessing personalized access
We recover the policies expected for CIFAR-$3A$.
By design, the CIFAR-$3A$ task compels “varying” profiles: Modiste’s forte. We find that Modiste learns to reconstruct near-optimal policies in both the standard and cost-aware settings, as depicted in Figure 4, for unseen users. Modiste learns online to provide users with model predictions for two classes, with expert consensus for another two classes, and no support for the final class. This is reflected quantitatively in Table 3, where both Modiste variants, particularly KNN, have lower expected losses than any of the offline policies. Using our human-informed synthetic decision-makers, we also identify a choice of $\lambda$ for each algorithm that modulates loss versus cost, which leads to a policy that uses the cheaper form of support (H-Only) more. However, the trade-off parameter affects the method differently, as the gap between the KNN variants is much smaller than that between the LinUCB variants.
Figure 5: Snapshots of the learned decision support policies computed at the end of $60$ interactions for MMLU-$3A$. The forms of support are colored in t-SNE embedding space. All six participants achieve $L_{h}(\pi)=0.1875$—comparable to the loss achieved by H-LLM—but exhibit distinct policies across input space. Modiste learns policies that capture each decision-maker’s strengths and weaknesses. All other participant plots are in Appendix Figure 11.
We learn to provide LLM access selectively for MMLU-$2A$.
Polymaths are rare; we observe no different in our Prolific data, as most participants struggle with at least one of the MMLU topics. As a result, Modiste learns that many individuals benefit from receiving the LLM as support, due in part to our task design (wherein the LLM excels at three of the four topics). In Figure 5, we see that Modiste personalizes support to decision-maker expertise, yielding policies that provide support on different topics for different decision-makers. Some users benefit from having the LLM all of the time (“strictly better” profile) while others can do without LLM for some topics. In Table 3, Modiste is not statistically different from the best baselines (H-LLM) in terms of expected loss. Yet, we observe lower expected costs for both LinUCB and KNN in the cost-agnostic setting, and find similar results, as expected, in the cost-aware setting. This suggests that learning policies with Modiste has merit: participants who have no expertise can expect to receive support from the LLM, and those who are proficient in one topic, but not in another, can expect to see the LLM in only cases beyond their expertise. While $L_{h}(\pi)$ of the Modiste variants are similar to that of the Population baseline in Table 3, the variance is significantly smaller—particularly with KNN—underscoring the benefits of personalization.
Discussion
Our human subject experiments validate trends observed in the computational experiments in both the cost-agnostic and cost-aware settings, demonstrating that Modiste can be used to learn decision support policies online. We find that learning personalized policies leads to performance and cost benefits when encountering users with “varying” expertise profiles, and recovers fixed policies for participants with “strictly better” profiles in the cost-agnostic setting. While our human-informed synthetic decision-makers permit us to pick a suitable $\lambda$ for unseen study participants, future work modeling human idiosyncrasies (Steyvers and Kumar 2022) explicitly in synthetic decision-makers may bolster the effectiveness of personalization.
Practical Considerations
Conclusion
A decision support policy captures when and which form of support should be provided to improve a decision-maker’s performance. The selective use of AI-based decision support helps instantiate the “appropriate use” clauses in emerging regulation (Biden 2023), as we only provide AI assistance to decision-makers as and when it is beneficial to them. We introduce Modiste, an interactive tool for learning a decision support policy for each decision-maker using contextual bandits. To the best of our knowledge, we are the first to learn and validate such a policy online for unseen decision-makers. Within our Modiste interface, we instantiate two variants of Algorithm 1 using existing stochastic contextual bandit tools, namely LinUCB and online KNN. We consider both cost-agnostic and cost-aware policies for providing decision-makers with access to decision support. Our computational and human subject experiments highlight the importance—and feasibility—of personalizing decision support policies for individual decision-makers. Our human subject experiments are encouraging, demonstrating that we can learn decision support policies in remarkably few iterations and tease apart differences in decision-makers’ need for support. While encouraging rich cross-talk between domain experts and practitioners, future work integrating Modiste into existing decision-making workflows would pave a route towards responsible use of AI as decision support.
@@ Conclusion
While we obtain promising results with Modiste, we acknowledge that we only consider the classification setting here where we get immediate feedback (e.g., we can calculate the loss to update $\pi$). We imagine that extending to a delayed feedback setting or to a different cognitive task (e.g., planning or perception) would prove fruitful. As with most online learning algorithms, selecting a suitable explore-exploit trade-off (a la Table 6) can prove challenging, though we find that our learned policy is preferable to the fixed policies used in practice at the moment. Though Modiste is promising, we note that significant issues can arise when decision-makers blindly rely on decision support (Buçinca et al. 2020), especially when the support is erroneous or ineffective; such over-reliance requires careful attention to prevent. Second, our problem definition hinges on domain experts defining the available forms of support (i.e., we need a clearly defined $\mathcal{A}$ to use Modiste). In practice, this may prove difficult, as one may not know how to define specific forms of support or decision-makers may have access to varying support sets due to regulatory or organizational reasons. Nonetheless, the promise of using ML models as decision support rests on our ability to provide the right form of support to the right decision-maker at the right time: Modiste takes a step towards realizing personalized access to decision support.
@@ Acknowledgements
UB acknowledges support from DeepMind and the Leverhulme Trust via the Leverhulme Centre for the Future of Intelligence (CFI). AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1, EPSRC grants EP/V056522/1 and EP/V056883/1, and the Leverhulme Trust via CFI. KMC acknowledges support from a Marshall Scholarship and the Cambridge Trust. PK acknowledges support from The Alan Turing Institute. AT acknowledges support from the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, an Amazon Web Services Award, a Facebook Faculty Research Award, funding from Booz Allen Hamilton Inc., and a Block Center Grant. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies. We thank the Prolific participants who took part in our studies, as well as our Cambridge-based volunteer pilot participants. We thank Alan Liu and Kartik Chandra for their invaluable input on designing our web-based server.
Appendix A Appendix
We provide details on decision support policies, additional computational experiments with Modiste, and extensive information on our human subject experiments with Modiste.
Appendix B Comparison Against Prior Work

### v3

@@ 6 Computational Evaluation
To construct expertise profiles with realistic values for each $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$, we collect data on user decisions across different users and then calculate individual $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$. The set of participant expertise profiles form a population of decision-makers that we refer to as human-informed synthetic decision-makers. From the estimated $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$ of each human-informed synthetic decision-maker, we can simulate decision-maker behavior.
To construct human-informed synthetic decision-makers, we recruited 20 participants from Prolific (10 for CIFAR-$3A$ and 10 for MMLU-$2A$). We use the same recruitment scheme as the larger human subject experiment described in the Appendix. We define regions of expertise over class labels for CIFAR-$3A$ and over question topics for MMLU-$2A$, as we expect $r_{\textsc{A}_{i}}(x;h)$ to be roughly constant for $x\in{\mathcal{X}}_{j}$ where ${\mathcal{X}}_{j}$ is defined by a class label or question topic. We showed each participant similar inputs with different forms of support to estimate $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$ for each support $\textsc{A}_{i}$ in each region ${\mathcal{X}}_{j}$. On each trial, each participant is randomly assigned a form of support; trials are approximately balanced by the type of support and grouping (i.e., topic or class). We compute participant accuracy averaged over all trials: 100 for CIFAR-$3A$, 60 for MMLU-$2A$. We denote expertise profiles as follows: if there were three regions in the input space, an individual’s expertise profile under support $\textsc{A}_{i}$ would be written as $r_{\textsc{A}_{i}}=[0.7,0.1,0.7]$, meaning the individual incurs a loss of $0.7$ on ${\mathcal{X}}_{1}$, $0.1$ on ${\mathcal{X}}_{2}$, and $0.7$ on ${\mathcal{X}}_{3}$.
@@
  • Approximately Invariant expertise across all the regions under different forms of support, i.e., $r_{\textsc{A}_{1}}({\mathcal{X}}_{j};h)\approx r_{\textsc{A}_{2}}({\mathcal{X}}_{j};h)\approx\cdots\approx r_{\textsc{A}_{k}}({\mathcal{X}}_{j};h),\forall j\in[N]$. Both a random decision support policy and a policy that selects a fixed form of support would suffice for such a profile.
@@
We investigate how personalized policies compare against offline baselines under each expertise profile (Table 1). We verify that learning decision support policies are not helpful for decision-makers with “invariant” expertise profiles. For individuals who fall under the “varying” profiles, we find at least one personalized policy outperforms offline policies and learns a policy that is significantly closer to the decision-maker’s optimal performance. This is because a personalized policy identifies which form of support is better in each context, compared to fixed offline policies which always show one form of support or to the population-wide variant, which may not provide the correct form of support to each individual. For individuals in the “strictly better” profile, while we do not find a statistically significant difference from the baselines due to the large variance in fixed policies (e.g., they work well for some decision-makers but poorly for others), we observe that much smaller variance with Modiste, particularly using KNN. For most individuals, the population-wide baseline performs poorly, emphasizing the need for personalization of decision support. Misalignment between the population policy and the optimal policy of the new decision-maker demonstrably leads to an ineffective use of decision support. We note that KNN generally outperforms LinUCB, the latter of which can be saddled by its implicit linearity assumption. We further study the effect of various parameters, e.g., exploration parameters, KNN parameters, embedding size, and the number of interactions in the Appendix.
7 Modiste with Real Users
@@
To validate whether Modiste can improve decision-maker performance in practice, we run a series of human subject experiments (i.e., ethics-reviewed studies with real human participants). We first introduce the set-up of our user study; additional information can be found in the Appendix.
@@
We recruit a total of $80$ crowdsourced participants from Prolific to interact with Modiste ($N=30$ and $N=50$ for CIFAR-$3A$ and MMLU-$2A$, respectively). We recruit more participants for MMLU-$2A$, as we expect greater individual differences in regions where support is needed, e.g., some participants may be good at mathematics and struggle in biology, whereas others may excel in biology.
@@
Each participant is assigned to only one task. Within a task, participants are randomly assigned to one algorithm variant; an equal number of participants are included per variant (i.e., 10 for MMLU and 5 for CIFAR). Participants are required to reside in the United States and speak English as a first language. Participants are paid at a base rate of \$9/hr, and are told they may be paid an optional bonus up to \$10/hr based on the number of correct responses. We allot 25-30 minutes for the CIFAR task and 30-40 for MMLU, as each MMLU question takes more effort. We applied the bonus to all participants in all studies. We run an ANOVA with Tukey HSD across the conditions for each task.
Modiste outperforms baselines for “varying” expertise profiles.
By design, the CIFAR-$3A$ task compels “varying” profiles: Modiste’s forte. This is reflected quantitatively in Figure 4 (left), where both Modiste variants have lower expected losses than any of the offline policies. In particular, KNN achieves statistically significantly lower expected loss over all variants ($p<0.001$ across all pairs), including the population-wide policy learned from the pilot study. When visualizing the learned decision support policies, we observe that Modiste can reconstruct near-optimal policies, as depicted in the Appendix.
Figure 4: We report expected average loss $L_{h}(\pi)$ (lower is better) and standard error in the last 10 trials by Prolific participants for each algorithm, with CIFAR conditions on the left and MMLU conditions on the right. In the CIFAR setting, where individuals typically exhibit “varying” expertise profiles, we see significant benefits from using Modiste, particularly in the KNN setting. While we observe that most individuals in the MMLU condition exhibit “strictly better” expertise, which means personalized policies typically only perform as well as the best baseline, we still observe instances of deferred decisions to the human on a case-by-case basis—see Figure 5.
Figure 5: Snapshots of the learned decision support policies computed at the end of the study for 10 participants on the MMLU task. The forms of support are colored in t-SNE embedding space. All participants exhibit distinct policies across input space. The bar plot to the right of each scatter plot shows the relative performance of that decision-maker alone in each category, ordered from left to right as M=Mathematics, B=Biology, CS=Computer Science, FP=Foreign Policy per subplot. When a decision-maker performs well alone, Modiste learns policies to empower that decision-maker without LLM access. For example, the individual in the top left is highly competent at both Mathematics and Foreign Policy; the learned decision support policy reflects this.
Modiste matches the best baseline for “strictly better” profiles.
Polymaths are rare; we observe no different in our human subject experiments on MMLU, as most participants are better with LLM access, which places them in the “strictly better” category. While participants with both Modiste variants outperform the H-Only baseline ($p<0.01$), we observe that on average Modiste settings are no different than H-LLM, the condition where people always have access to the LLM, as shown in Figure 4 (right). This confirms our hypothesis from the computational experiments. Further, while the average performance of Modiste variants is similar to that of the offline fixed or population-wide policies in Figure 4 (right), the variance is significantly smaller—particularly with KNN.
Modiste can facilitate appropriate use of decision support.
Since the LLM only excels at three of the four topics on the MMLU task, Modiste learns in many cases to defer to human judgment on Mathematics questions, particularly when the human has strong expertise. In Figure 5, we visualize the learned decision support policies of various individuals in the study and illustrate how Modiste yields policies that provide support on different topics for different decision-makers.
Limitations.
@@
In this work, we consider the classification setting where we get immediate feedback (e.g., we can calculate the loss to update $\pi$). Future work can consider more complex decision-making tasks that may require extending to a delayed feedback setting or to a different cognitive task (e.g., planning or perception). Though Modiste is promising, we note that significant issues can arise when decision-makers blindly rely on decision support (Buçinca et al. 2020; Chen et al. 2023), especially when the support is erroneous or ineffective; such over-reliance requires careful attention to prevent. Further, our problem definition hinges on domain experts defining the available forms of support (i.e., we need a clearly defined $\mathcal{A}$ to use Modiste). In practice, this may prove difficult, as one may not know how to define specific forms of support or decision-makers may have access to varying support sets.
8 Conclusion
A decision support policy captures when and which form of support should be provided to improve a decision-maker’s performance. The selective use of AI-based decision support helps instantiate the “appropriate use” clauses in emerging regulation (Biden 2023), as we only provide AI assistance to decision-makers as and when it is beneficial to them. We introduce Modiste, an interactive tool for learning a decision support policy for each decision-maker using contextual bandits. To the best of our knowledge, we are the first to learn and validate such a policy online for unseen decision-makers. Within our Modiste interface, we instantiate two variants of Algorithm 1 using existing stochastic contextual bandit tools, namely LinUCB and online KNN. Our computational and human subject experiments highlight the importance—and feasibility—of personalizing decision support policies for individual decision-makers. Our human subject experiments show promise, as we personalize decision support policies in few iterations yet find nuances in decision-makers’ need for support: some unskilled decision-makers uniformly benefit from LLM access, while others only need LLMs for some tasks. While encouraging rich cross-talk between domain experts and practitioners, future work integrating Modiste into existing decision-making workflows would pave a route towards responsible use of AI as decision support.
@@ Acknowledgements
UB acknowledges support from DeepMind and the Leverhulme Trust via the Leverhulme Centre for the Future of Intelligence (CFI). AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1, EPSRC grants EP/V056522/1 and EP/V056883/1, and the Leverhulme Trust via CFI. KMC acknowledges support from a Marshall Scholarship and the Cambridge Trust. PK acknowledges support from The Alan Turing Institute. AT acknowledges support from the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, an Amazon Web Services Award, a Facebook Faculty Research Award, funding from Booz Allen Hamilton Inc., and a Block Center Grant. This work is supported (in part) by ELSA - European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of any of these funding agencies, European Union or European Commission. We thank the Prolific participants and our volunteer pilot participants. We thank Alan Liu and Kartik Chandra for their invaluable input on designing our web-based server.
Appendix A Comparison Against Prior Work